OpenStudy (lena772):

If f(x) is the total cost, in dollars, of x candies, which of the following statements best describes the meaning of f(2) = 6? The total cost of 2 candies is $6.00. The total cost of 6 candies is $2.00. The total cost of 2 candies is $3.00. The total cost of 3 candies is $2.00.

OpenStudy (anonymous):

think of f(x) as "y" so f(x) = x^2 for example is simply y = x^2

OpenStudy (anonymous):

Now in f(x) x is the quantity for which x = #

OpenStudy (lena772):

The first one?

OpenStudy (anonymous):

so if we had: f(x) = 3x this meant that 3 multiplied by a number (x) equals y or f(x) f(x) = 3x if we had 2, then it would be x = 2 so: f(2) = 3(2) f(2) = 6

OpenStudy (lena772):

Thanks

OpenStudy (anonymous):

yw (: